# A Survey on Resource Management in Joint Communication and Computing-Embedded SAGIN

### v1

@@ II-C Evolution of Network Integration in JCC-SAGIN
Recent advancements in network integration, reflecting contributions from both academia and industry, are summarized in Table VIII. International organizations such as 3rd Generation Partnership Project (3GPP) and European Telecommunications Standards Institute (ETSI) have proposed standards for integrating various SAGIN segments, extending from physical to application layers [46, 47, 48, 49]. Additionally, entities like China Unicom [50] and China Future Mobile Communication Forum [51] have published white papers highlighting the characteristics of network components, typical applications, and the impact of technology on SAGIN. Beyond academic research, several corporations have initiated trials in network integration. For instance, EdgeSAT [52] and Viasat [53] have concentrated on integrating satellite and TN, while Facebook [54] and Nokia [55] have developed air-terrestrial integration networks (ATIN).
@@
TABLE VIII: Academic and Industrial Trials on the network integration
@@
ETSI [49] 2023-05 STIN Detail the radio resource management requirements for frequency division duplexing (FDD) and time division duplexing (TDD) modes in E-UTRA, including measurement requirements for UTRAN and users, and the demand for dynamic node behavior and interaction, such as instant latency and response characteristics.
@@
Industry Trial EdgeSAT [52] 2022-03 STIN Illustrate the potential evolution of SN to address specific use cases, focusing on seamless integration with terrestrial access networks and backhaul links, and managing resources of edge nodes to enhance scalability, especially in scenarios involving numerous edge nodes.
@@ III-C3 RIS-Assisted Cooperative MEC framework in JCC-SAGIN
Key issues in aerial RIS-assisted MEC include resource management, trajectory planning, and beamforming design. Shang et al. [96] have highlighted these challenges, along with the need to balance excessive path loss in Non-Line-of-Sight (NLoS) conditions. To fully exploit the advantages of MEC and RIS, a max-min computation capacity problem is formulated under the Orthogonal Multiple Access (OMA) scheme [97]. The simulation results verified that the proposed scheme could achieve an 8.08 Mb computation capacity higher than the case without RIS. The authors extended their model and discussed the computation capacity maximization problem in the RIS-aided UAV-MEC system with Non-Orthogonal Multiple Access (NOMA). The simulation results showed that compared to the OMA scheme, the NOMA scheme could achieve higher computation capacity [98]. Zhai et al. [99] and Mei et al. [100] investigated the EE issues in RIS-assisted UAV networks.
@@ III-D Digital Twin-Enabled JCC-SAGIN
Existing works on DT mainly focused on combining it with MEC and AI technology. Aided by DT, the offloading unit does not always need to interact with the physical environment and store the real-time status of each platform. Therefore, compared to the MEC network without DT, the system performance and user experience in the DT edge network can improve significantly [104]. Up to now, the research on DT edge networks in SAGIN is still in the early stages. DT is first used to train the proactive DRL scheme offline in a UAV-assisted MEC system to tackle the asynchronous training of deep recurrent Q-network (DRQN) [105]. To capture the time-varying features of SAGIN, dynamic DT is integrated with FL to maximize the clients’ utilities, where the DT of the UAV acts as the leader of the formulated Stackelberg game [106]. This is the first work to consider the unpredictable and mobile network topology during the FL procedure in the air-ground network. Different from previous works, which assume that the mobility and task requirements can be predetermined, Sun et al. proposed a DT-driven two-stage incentive scheme to jointly maximize the users’ satisfaction and overall EE considering the limitations of UAVs [107]. Compared to the centralized resource allocation algorithm, the proposed decentralized method can decrease the computing burden of vehicles while ensuring system performance. The authors extended their research by considering the design of lightweight DT architecture in air-ground networks [108]. The integration of DT and FL with a distributed incentive mechanism is formulated to maximize learning efficiency.
@@ III-E Summary Remarks
This section examines the enabling technologies driving improvements in network communication and computing in JCC-SAGIN, focusing on network slicing, EH, RIS, and DT technologies. SAGIN’s ability to cater to a wide variety of service demands overcomes the limitations of traditional physical network architectures in delivering on-demand services for diverse scenarios. Network slicing emerges as a vital solution, creating multiple virtual logical subnetworks within the same physical network infrastructure, thus addressing the need for customizable and flexible network services. EH and WPT are pivotal in promoting sustainable network operations, enabling devices to collect energy from their environment or the network platforms, thereby alleviating the need for frequent recharging and battery replacements. Moreover, the application of RIS is shown to significantly enhance future communication networks by supporting computing task offloading and improving network computing efficiency, alongside boosting communication capabilities. However, the implementation of RIS in satellites and UAVs introduces challenges associated with the need for regular configuration adjustments. Furthermore, the adoption of DT technology in SAGIN facilitates intelligent decision-making and effective resource management, significantly enhancing network reliability and operational efficiency. This integration of enabling technologies underscores a strategic approach to overcoming the inherent challenges of SAGIN, setting a foundation for the next generation of network communication and computing performance.
@@ IV-G Summary Remarks
This section underscores the wide range of application scenarios enabled by JCC-SAGIN, alongside listing engineering examples that illustrate the network architecture’s adaptability. Within the diverse use cases highlighted, components such as satellites and UAVs are tasked with not just fulfilling basic communication needs but also delivering computational services for complex and computation-intensive tasks like remote sensing data processing, image recognition, and machine learning (ML) applications. This dual requirement emphasizes the necessity and critical importance of addressing joint communication and computation challenges in SAGIN. Effective resource management in JCC-SAGIN plays a crucial role in enhancing system performance, minimizing costs, and aligning with user demands more effectively. Resource management covers various aspects, such as user association policies, allocation of power and computing resources, beamforming, and UAV trajectory optimization. The strategic management of these resources is pivotal for the design and operation of future 6G SAGIN, a topic that is thoroughly explored in Section VI. This focus on resource management is integral to unlocking the full potential of JCC-SAGIN, ensuring it meets the evolving demands in the future.
@@ V-A1 Task Completion Time
In JCC-SAGIN, task completion time encompasses both the transmission time for communication and the computing time for executing tasks. Transmission time is influenced by the allocation of communication resources (like bandwidth and power) and the distance of communication, whereas processing time depends on the allocation of computing resources and task offloading strategies. Zhang et al. examined the JCC issue in Satellite Communication Networks (SCNs) and introduced a game theory-based approach to minimize total weighted delay for computationally intensive tasks [121]. In another study, the authors considered the interplay between communication and computation, developing a broadband aggregation algorithm for over-the-air federated learning (AirFL) to reduce communication delays [122]. Previous research in JCC-TN often assumes the computed results entail minimal data, overlooking the return time of these results. However, in JCC-SAGIN, the considerable propagation delay between satellites and the ground necessitates accounting for return time, highlighting a distinct difference in JCC problem research between SAGIN and TN.
@@ V-C1 Platform Distribution
While the PPP model yields accurate results for TN performance analysis, it is less applicable to communication scenarios involving a limited number of SAPs dispersed over a given area [160]. The Binomial point process (BPP) is a more suitable geometric method for modeling the spatial distribution of a finite number of SAPs [161]. Table IX presents several notable studies on the performance analysis of NTN. Chetlur et al. investigated the coverage probability of downlink UAV networks, with UAVs modeled as a 3D BPP and assuming UAV-to-user links undergoing Nakagami-m fading [162]. Likewise, research on satellite networks has modeled the location distribution of LEO satellites as a 3D BPP [163, 164]. Okati et al. explored satellite network performance under a nonhomogeneous PPP (NPPP) model, analyzing communication variances for users located at different latitudes [165, 166]. This approach extends the understanding of NTN performance by considering the distinct spatial distributions and characteristics of SAPs, as opposed to traditional TN.
TABLE IX: Typical Related works of performance analysis for NTN
@@ V-C2 Performance Analysis
* Numerical and Theoretical Analysis: Coverage in SAGIN is typically measured by signal outage probability, defined as the likelihood of received SNR or SINR exceeding a threshold. Capacity analysis often relies on ergodic capacity derived from Shannon’s formula. Shadowed-Rician (SR) fading has been identified as a more appropriate model for small-scale fading and shadowing in SAP-to-ground channels compared to traditional Nakagami-m fading [175]. Huang et al. formulated a constrained optimization problem to maximize system capacity, derived closed-form expressions for the outage probability and ergodic capacity in hybrid satellite-terrestrial systems under SR fading [176], and proposed two multi-user scheduling schemes (best user scheduling and user fairness scheduling) to improve system performance. Bhatnagar et al. investigated the maximum ratio combining (MRC) scheme in the land mobile satellite (LMS) channel under SR fading channels [177]. This study derived approximate closed-form expressions for the bit error rate, outage probability, and capacity of the considered scheme. The resulting bit error rate expression is particularly valuable for assessing the coding gain of the MRC schem. However, system-level analysis for SR fading channels is limited. Na et al. and Jung et al. analyzed outage probability and coverage performance in satellite networks under SR fading, but did not fully address co-channel interference [167, 163].
@@ V-E Summary Remarks
In this section, we reviewed resource management modeling in JCC-SAGIN from four aspects: metrics, offloading strategies, network modeling, and service modeling. Table X summarizes the typical related works on offloading in JCC-SAGIN and analyzes them concerning scenario, computation framework, computation strategy, and optimization objective.
@@
TABLE X: Summary of Related works on offloading in JCC-SAGIN
@@ VI-A1 User-Platform Association
Common integer programming methods are summarized in Table XI [208]. Let’s denote the association variable between user $k$ and SAP $i$ as $\alpha_{k,i}\in\left\{0,1\right\}$. Continuous methods typically handle binary constraints in two ways. The first is through relaxation-based methods, where binary constraints are relaxed into a continuous range, $\alpha_{k,i}\in\left[0,1\right]$. These can be further categorized into approximate and exact methods. Approximate methods round continuous solutions to discrete values. For instance, Yu et al. used a relaxation-based (approximate) method in discussing UAV-BS association optimization [148], omitting one constraint for easier subproblem reconstruction before recovering binary solutions. However, this approach doesn’t guarantee local optimality due to separate optimization and rounding processes. Exact methods like Branch-and-Bound and cutting plane approaches avoid rounding but require repetitive linear relaxations, leading to longer runtimes.
The second continuous method is replacement-based, substituting the concerned space with a set of equivalent continuous variables. This often involves designing penalty functions and solving non-convex optimization problems iteratively, a process that can be complex and time-consuming. Wu et al. proposed the $\ell_{p}$-Box ADMM algorithm to bridge integer programming and continuous optimization, enhancing convergence speed and computational efficiency [208].
TABLE XI: Summary of Integer Programming Methods
@@ VI-B2 Deep Learning
Deep neural networks (DNNs) are often applied for augmented reality applications, natural language processing, and computer vision [243]. Based on DNN, a two-layer optimization algorithm is presented to solve the nonlinear programming problem in a multi-UAV-enabled MEC system [111]. The upper layer algorithm utilizes differential evolution (DE) to obtain the optimal location of UAVs, while the lower algorithm uses distributed DNN (DDNN) to address the offloading problem. Compared to a single DNN, the adopted multiple DNNs accelerate the convergence speed and lead to significant differences in the outer layer.
@@ VI-B3 Reinforcement Learning
Other learning algorithms are also developed based on DQN. A dueling DQN is proposed to determine the node selection and block size decision in the UAV-assisted MEC system, which can calculate the state value and action reward. The adopted dueling DQN can also increase the estimating capability of the environmental state to achieve the optimization objective [251]. Since the DQN algorithm may result in overestimating the Q value, double DQN (DDQN) can overcome this disadvantage by separating the selection action from the evaluation one. In order to minimize energy consumption, a DDQN algorithm is proposed to jointly optimize the user association and UAV trajectory in the DT-assisted UAV-enabled MEC system [104]. An online DDQN framework is also adopted in air-ground integrated edge computing networks, where the two DQNs are used to generate the Q-factor and the post-decision Q-factor of the mobile device [252]. However, the sampling from reply memory is random in the conventional double DQL, neglecting the timeliness and heterogeneity of different actions. This fact is improper for delay-sensitive applications. Aiming at improving the sampling efficiency, a latency-sensitive replay memory algorithm based on double DQL is presented to obtain the offloading strategy by utilizing the local and neighboring historical information [253]. Furthermore, a deep reinforcement learning (DRL) offloading method based on dueling double DQN (D3QN) addresses the MDP formulation. By combining the DDQN and the dueling framework, the D3QN-based offloading policy trains the approximating function and calculates the Q value of each action at each state [144].
@@
* Summary Remarks: Plenty of works investigated the applications of RL methods in resource management due to its advantage of model free. Table XII summarizes and compares the typical works which introduces RL method to provide intelligent and adaptive solutions for resource management in JCC-SAGIN. Most works assume that there is a server for centralized training, which may lead to higher costs in SAGIN than in TN. In addition, the training and execution process will introduce additional latency and energy consumption. These facts will incur little performance loss on delay and increase the burden of energy-limited devices. Therefore, it is necessary to design adpative DTDE algorithms, which consider different features of platforms or devices and their deployed resources in JCC-SAGIN.
TABLE XII: Typical works using Reinforcement Learning Method in JCC-SAGIN
@@
[253] SAGIN Delay-sensitive replay memory algorithm (DSRPM) High dynamic and complex optimization issues Reduce package loss and increase throughput, improving packet delay

### v3

@@ II-C Evolution of Network Integration in JCC-SAGIN
Recent advancements in network integration, reflecting contributions from both academia and industry, are summarized in Table IX. International organizations such as 3rd Generation Partnership Project (3GPP) and European Telecommunications Standards Institute (ETSI) have proposed standards for integrating various SAGIN segments, extending from physical to application layers [46, 47, 48, 49]. Additionally, entities like China Unicom [50] and China Future Mobile Communication Forum [51] have published white papers highlighting the characteristics of network components, typical applications, and the impact of technology on SAGIN. Beyond academic research, several corporations have initiated trials in network integration. For instance, EdgeSAT [52] and Viasat [53] have concentrated on integrating satellite and TN, while Facebook [54] and Nokia [55] have developed air-terrestrial integration networks (ATIN).
@@
TABLE IX: Academic and Industrial Trials on the network integration
@@
ETSI [49] 2023-05 STIN Detail the radio resource management requirements for frequency division duplexing (FDD) and TDD modes in E-UTRA, including measurement requirements for UTRAN and users, and the demand for dynamic node behavior and interaction, such as instant latency and response characteristics.
@@
Industry Trial EdgeSAT [52] 2022-03 STIN Illustrate the potential evolution of satellite networks to address specific use cases, focusing on seamless integration with terrestrial access networks and backhaul links, and managing resources of edge nodes to enhance scalability, especially in scenarios involving numerous edge nodes.
@@ III-C3 RIS-Assisted Cooperative MEC framework in JCC-SAGIN
Key issues in aerial RIS-assisted MEC include resource management, trajectory planning, and beamforming design. Shang et al. [96] have highlighted these challenges, along with the need to balance excessive path loss in non-LoS (NLoS) conditions. To fully exploit the advantages of MEC and RIS, a max-min computation capacity problem is formulated under the orthogonal multiple access (OMA) scheme [97]. The simulation results verified that the proposed scheme could achieve an 8.08 Mb computation capacity higher than the case without RIS. The authors extended their model and discussed the computation capacity maximization problem in the RIS-aided UAV-MEC system with non-orthogonal multiple access (NOMA). The simulation results showed that compared to the OMA scheme, the NOMA scheme could achieve higher computation capacity [98]. Zhai et al. [99] and Mei et al. [100] investigated the EE issues in RIS-assisted UAV networks.
@@ III-D Digital Twin-Enabled JCC-SAGIN
Existing works on DT mainly focused on combining it with MEC and AI technology. Aided by DT, the offloading unit does not always need to interact with the physical environment and store the real-time status of each platform. Therefore, compared to the MEC network without DT, the system performance and user experience in the DT edge network can improve significantly [104]. Up to now, the research on DT edge networks in SAGIN is still in the early stages. DT is first used to train the proactive deep reinforcement learning (DRL) scheme offline in a UAV-assisted MEC system to tackle the asynchronous training of deep recurrent Q-network (DRQN) [105]. To capture the time-varying features of SAGIN, dynamic DT is integrated with FL to maximize the clients’ utilities, where the DT of the UAV acts as the leader of the formulated Stackelberg game [106]. This is the first work to consider the unpredictable and mobile network topology during the FL procedure in the air-ground network. Different from previous works, which assume that the mobility and task requirements can be predetermined, Sun et al. proposed a DT-driven two-stage incentive scheme to jointly maximize the users’ satisfaction and overall EE considering the limitations of UAVs [107]. Compared to the centralized resource allocation algorithm, the proposed decentralized method can decrease the computing burden of vehicles while ensuring system performance. The authors extended their research by considering the design of lightweight DT architecture in air-ground networks [108]. The integration of DT and FL with a distributed incentive mechanism is formulated to maximize learning efficiency.
@@ III-E Summary Remarks
This section examines the enabling technologies driving improvements in network communication and computing in JCC-SAGIN, focusing on network slicing, EH, RIS, and DT technologies. Table X summarizes the capabilities of these enabling technologies. SAGIN’s ability to cater to a wide variety of service demands overcomes the limitations of traditional physical network architectures in delivering on-demand services for diverse scenarios. Network slicing emerges as a vital solution, creating multiple virtual logical subnetworks within the same physical network infrastructure, thus addressing the need for customizable and flexible network services. EH and WPT are pivotal in promoting sustainable network operations, enabling devices to collect energy from their environment or the network platforms, thereby alleviating the need for frequent recharging and battery replacements. Moreover, applying RIS significantly enhances future communication networks by supporting computing task offloading, improving network computing efficiency, and boosting communication capabilities. Furthermore, adopting DT technology in SAGIN facilitates intelligent decision-making and effective resource management, significantly enhancing network reliability and operational efficiency. The integration of enabling technologies underscores a strategic approach to overcoming SAGIN’s inherent challenges, setting a foundation for the next generation of network communication and computing performance.
TABLE X: Capabilities of the Enabling Technologies in JCC-SAGIN
Enabling technologies Ability and contribution to improve the performance of the JCC-SAGIN
Network Slicing, SDN, and NFV Provide on-demand services, facilitate swift deployment and adaptable adjustment capabilities, enhance resource utilization efficiency and resource management efficacy, and achieve heightened flexibility and scalability.
EH and WPT Ensure continuous communication and computing, enhance network sustainability, facilitate more flexible deployment and layout, and reduce energy consumption and costs.
RIS Assist in communication and computing task offloading, facilitate signal and coverage enhancement, improve spectral and energy efficiency. Offer high flexibility, adaptability, and reconfigurability.
DT Beneficial for achieving efficient resource management and scheduling, intelligent decision-making, fault diagnosis, and risk assessment, thereby enhancing network reliability and stability.
@@ IV-F2 Navigation for Search and Rescue Operations
TABLE XI: Enhancements in Sector Applications Through JCC-SAGIN and Resource Management
Applications Impact of JCC-SAGIN and resource management on various applications
Water monitoring JCC-SAGIN enables dynamic adjustments in water quality assessments and pollution control strategies, tailoring responses to environmental changes. Effective resource management optimizes water usage and conservation strategies, allowing for sustainable development and allocation of water resources across diverse ecosystems.
Smart oilfields By continuously monitoring conditions and automating responses, JCC-SAGIN minimizes downtime and maximizes oil production. Strategic resource management allows for precise allocation of drilling resources and maintenance efforts, ensuring operational efficiency while reducing environmental impact and operational costs.
Precision agriculture By creating a detailed view of crop health and soil conditions, JCC-SAGIN supports advanced predictive analytics for crop yields and resource use. Efficient resource management optimizes input usage like water and fertilizers, enhances land use efficiency, and reduces the ecological footprint of farming practices.
Wildlife conservation JCC-SAGIN boosts conservation efforts by providing detailed tracking of wildlife through satellite and ground-based monitoring networks. Resource management enable targeted conservation actions, optimal allocation of conservation funds, and quick mobilization of resources to protect threatened species and habitats.
Disaster management JCC-SAGIN streamlines communication between emergency teams and command centers, ensuring timely evacuations and effective deployment of resources. Through sophisticated resource management, JCC-SAGIN directs essential supplies and personnel efficiently, minimizing disaster impact and aiding in quicker recovery processes.
Search and rescue JCC-SAGIN reduces search times dramatically and improves survival rates by enabling precise and swift delivery of emergency aid. Advanced resource management enhances logistical coordination, ensuring that rescue operations are well-equipped and responders can access critical resources promptly during emergencies.
@@ IV-G Summary Remarks
This section underscores the wide range of application scenarios enabled by JCC-SAGIN, alongside listing engineering examples that illustrate the network architecture’s adaptability. The impact of JCC-SAGIN and resource management on various applications have been listed in Table XI. Within the diverse use cases highlighted, components such as satellites and UAVs are tasked with not just fulfilling basic communication needs but also delivering computational services for complex and computation-intensive tasks like remote sensing data processing, image recognition, and ML applications. This dual requirement emphasizes the necessity and critical importance of addressing joint communication and computation challenges in SAGIN. Effective resource management in JCC-SAGIN plays a crucial role in enhancing system performance, minimizing costs, and aligning with user demands more effectively. Resource management covers various aspects, such as user association policies, allocation of power and computing resources, beamforming, and UAV trajectory optimization. The strategic management of these resources is pivotal for the design and operation of future 6G SAGIN, a topic that is thoroughly explored in Section VI. This focus on resource management is integral to unlocking the full potential of JCC-SAGIN, ensuring it meets the evolving demands in the future.
@@ V-A1 Task Completion Time
In JCC-SAGIN, task completion time encompasses both the transmission time for communication and the computing time for executing tasks. Transmission time is influenced by the allocation of communication resources (like bandwidth and power) and the distance of communication, whereas processing time depends on the allocation of computing resources and task offloading strategies. Zhang et al. examined the JCC issue in satellite communication networks and introduced a game theory-based approach to minimize total weighted delay for computationally intensive tasks [121]. In another study, the authors considered the interplay between communication and computation, developing a broadband aggregation algorithm for over-the-air federated learning (AirFL) to reduce communication delays [122]. Previous research in TN often assumes the computed results entail minimal data, overlooking the return time of these results. However, in JCC-SAGIN, the considerable propagation delay between satellites and the ground necessitates accounting for return time, highlighting a distinct difference in JCC problem research between SAGIN and TN.
@@ V-C1 Platform Distribution
While the PPP model yields accurate results for TN performance analysis, it is less applicable to communication scenarios involving a limited number of SAPs dispersed over a given area [160]. The Binomial point process (BPP) is a more suitable geometric method for modeling the spatial distribution of a finite number of SAPs [161]. Table XII presents several notable studies on the performance analysis of NTN. Chetlur et al. investigated the coverage probability of downlink UAV networks, with UAVs modeled as a 3D BPP and assuming UAV-to-user links undergoing Nakagami-m fading [162]. Likewise, research on satellite networks has modeled the location distribution of LEO satellites as a 3D BPP [163, 164]. Okati et al. explored satellite network performance under a nonhomogeneous PPP (NPPP) model, analyzing communication variances for users located at different latitudes [165, 166]. This approach extends the understanding of NTN performance by considering the distinct spatial distributions and characteristics of SAPs, as opposed to traditional TN.
TABLE XII: Typical Related works of performance analysis for NTN
@@ V-C2 Performance Analysis
* Numerical and Theoretical Analysis: Coverage in SAGIN is typically measured by signal outage probability, defined as the likelihood of received signal-to-noise ratio (SNR) or signal-to-interference-plus-noise ratio (SINR) exceeding a threshold. Capacity analysis often relies on ergodic capacity derived from Shannon’s formula. Shadowed-Rician (SR) fading has been identified as a more appropriate model for small-scale fading and shadowing in SAP-to-ground channels compared to traditional Nakagami-m fading [175]. Huang et al. formulated a constrained optimization problem to maximize system capacity, derived closed-form expressions for the outage probability and ergodic capacity in hybrid satellite-terrestrial systems under SR fading [176], and proposed two multi-user scheduling schemes (best user scheduling and user fairness scheduling) to improve system performance. Bhatnagar et al. investigated the maximum ratio combining (MRC) scheme in the land mobile satellite channel under SR fading channels [177]. This study derived approximate closed-form expressions for the bit error rate, outage probability, and capacity of the considered scheme. The resulting bit error rate expression is particularly valuable for assessing the coding gain of the MRC schem. However, system-level analysis for SR fading channels is limited. Na et al. and Jung et al. analyzed outage probability and coverage performance in satellite networks under SR fading, but did not fully address co-channel interference [167, 163].
@@ V-E Summary Remarks
In this section, we reviewed resource management modeling in JCC-SAGIN from four aspects: metrics, offloading strategies, network modeling, and service modeling. Table XIII summarizes the typical related works on offloading in JCC-SAGIN and analyzes them concerning scenario, computation framework, computation strategy, and optimization objective.
@@
TABLE XIII: Summary of Related works on offloading in JCC-SAGIN
@@ VI-A1 User-Platform Association
Common integer programming methods are summarized in Table XIV [208]. Let’s denote the association variable between user $k$ and SAP $i$ as $\alpha_{k,i}\in\left\{0,1\right\}$. Continuous methods typically handle binary constraints in two ways. The first is through relaxation-based methods, where binary constraints are relaxed into a continuous range, $\alpha_{k,i}\in\left[0,1\right]$. These can be further categorized into approximate and exact methods. Approximate methods round continuous solutions to discrete values. For instance, Yu et al. used a relaxation-based (approximate) method in discussing UAV-BS association optimization [148], omitting one constraint for easier subproblem reconstruction before recovering binary solutions. However, this approach doesn’t guarantee local optimality due to separate optimization and rounding processes. Exact methods like Branch-and-Bound and cutting plane approaches avoid rounding but require repetitive linear relaxations, leading to longer runtimes.
The second continuous method is replacement-based, substituting the concerned space with a set of equivalent continuous variables. This often involves designing penalty functions and solving non-convex optimization problems iteratively, a process that can be complex and time-consuming. Wu et al. proposed the $\ell_{p}$-Box alternating direction method of multipliers (ADMM) algorithm to bridge integer programming and continuous optimization, enhancing convergence speed and computational efficiency [208].
TABLE XIV: Summary of Integer Programming Methods
@@ VI-B2 Deep Learning
Deep neural networks (DNNs) are often applied for augmented reality applications, natural language processing, and computer vision [243]. Based on DNN, a two-layer optimization algorithm is presented to solve the nonlinear programming problem in a multi-UAV-enabled MEC system [111]. The upper layer algorithm utilizes differential evolution to obtain the optimal location of UAVs, while the lower algorithm uses distributed DNN (DDNN) to address the offloading problem. Compared to a single DNN, the adopted multiple DNNs accelerate the convergence speed and lead to significant differences in the outer layer.
@@ VI-B3 Reinforcement Learning
Other learning algorithms are also developed based on DQN. A dueling DQN is proposed to determine the node selection and block size decision in the UAV-assisted MEC system, which can calculate the state value and action reward. The adopted dueling DQN can also increase the estimating capability of the environmental state to achieve the optimization objective [251]. Since the DQN algorithm may result in overestimating the Q value, double DQN (DDQN) can overcome this disadvantage by separating the selection action from the evaluation one. In order to minimize energy consumption, a DDQN algorithm is proposed to jointly optimize the user association and UAV trajectory in the DT-assisted UAV-enabled MEC system [104]. An online DDQN framework is also adopted in air-ground integrated edge computing networks, where the two DQNs are used to generate the Q-factor and the post-decision Q-factor of the mobile device [252]. However, the sampling from reply memory is random in the conventional double DQL, neglecting the timeliness and heterogeneity of different actions. This fact is improper for delay-sensitive applications. Aiming at improving the sampling efficiency, a latency-sensitive replay memory algorithm based on double DQL is presented to obtain the offloading strategy by utilizing the local and neighboring historical information [253]. Furthermore, a DRL offloading method based on dueling double DQN (D3QN) addresses the MDP formulation. By combining the DDQN and the dueling framework, the D3QN-based offloading policy trains the approximating function and calculates the Q value of each action at each state [144].
@@
* Summary Remarks: Plenty of works investigated the applications of RL methods in resource management due to its advantage of model free. Table XV summarizes and compares the typical works which introduces RL method to provide intelligent and adaptive solutions for resource management in JCC-SAGIN. Most works assume that there is a server for centralized training, which may lead to higher costs in SAGIN than in TN. In addition, the training and execution process will introduce additional latency and energy consumption. These facts will incur little performance loss on delay and increase the burden of energy-limited devices. Therefore, it is necessary to design adpative DTDE algorithms, which consider different features of platforms or devices and their deployed resources in JCC-SAGIN.
TABLE XV: Typical works using Reinforcement Learning Method in JCC-SAGIN
@@
[253] SAGIN Delay-sensitive replay memory algorithm High dynamic and complex optimization issues Reduce package loss and increase throughput, improving packet delay